# Fabian Mauthe

### v2

@@ 86 UCI - Condition monitoring of hydraulic systems Data Set
ZeMA - Center for Mechatronics and Automation Technology Saarbrücken, Germany Mechatronic system
@@
Signals: efficiency, flow rate, power, pressure, temperature, vibration Diagnosis Link
@@ 87 UCI - Mechanical Analysis 1990
F. Bergadano, A. Giordana, L. Saitta, (1990). University of Torino, Italy Mechatronic system
@@
Signals: speed Diagnosis Link
@@ 88 UCI - Robot Execution Failures Data Set
Universidade Nova de Lisboa, Monte da Caparica, Portugal Robotic
@@
Signals: force, torque Diagnosis Link
@@ 89 UCI - Steel Plates Faults
Semeion Research Center of Sciences of Communication - Rome, Italy Materials
@@
Signals: geometry, inspection data, operating condition Diagnosis Link
@@ 90 Uni Lulea - Vibration Signals Wind Turbines
Lulea University of Technology: Martin del Campo Barraza, Sergio (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Sandin, Fredrik (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Strömbergsson, Daniel (Department of Engineering Sciences and Mathematics, Machine Elements) Mechanical component
@@
Signals: vibration Fault detection Link
@@ 91 Uni Paderborn KAt - Bearing Damage
University Paderborn Kat Mechanical component
@@
Signals: speed, temperature, torque Diagnosis Link
@@ 92 Virkler - Fatigue Crack Propagation
Virkler, D. A., Hillberry, B. M. and Goel, P. K. (1979). The Statistical Nature of Fatigue Crack Propagation. Journal of Engineering Materials and Technology, vol. 101, 148–153. Materials
@@
Signals: crack length Prognosis Link
@@ 93 WASEDA - Fault Detection and Classification
De Bruijn, B., Nguyen, T. A., Bucur, D., & Tei, K. (2016). Benchmark datasets for fault detection and classification in sensor data. In A. Ahrens, O. Postolache, & C. Benavente-Peces (Eds.), SENSORNETS 2016 - Proceedings of the 5th International Confererence on Sensor Networks, pp. 185-195. . SciTePress. https://doi.org/10.5220/0005637901850195 Electrical component
@@
Signals: light, temperature Diagnosis Link
@@ 94 Zenodo - Ball bearings subjected to time-varying load and speed conditions
Aimiyekagbon, O. K. (2024). Run-tofailure data set of ball bearings subjected to time-varying load and speed conditions. Zenodo, doi: 10.5281/zenodo.10805043, University Paderborn Mechanical component
@@
Signals: temperature, vibration Prognosis Link
@@ 95 Zenodo - Data-driven capacity estimation of commercial lithium-ion batteries from voltage relaxation
Jiangong Zhu, Yixiu Wang, Yuan Huang, R. Bhushan Gopaluni, Yankai Cao, Michael Heere, Martin J. Mühlbauer, Liuda Mereacre, Haifeng Dai, Xinhua Liu, Anatoliy Senyshyn, Xuezhe Wei, Michael Knapp, & Helmut Ehrenberg. (2022). Data-driven capacity estimation of commercial lithium-ion batteries from voltage relaxation [Data set]. Zenodo. https://doi.org/10.5281/zenodo.6405084 Electrical component
@@
Signals: current, temperature, voltage Prognosis Link
@@
Zhang, Y., Tang, Q., Zhang, Y., Wang, J., Stimming, U., & Lee, A. A. (2020). Identifying degradation patterns of lithium ion batteries from impedance spectroscopy using machine learning [Data set]. Zenodo. https://doi.org/10.5281/zenodo.3633835 Electrical component
@@
Signals: current, temperature, voltage Prognosis Link
@@ 97 Zenodo - Predictive maintenance dataset
Huawei Munich Research Center Mechatronic system
@@
Signals: humidity, vibration Prognosis Link
@@ 98 Zenodo - TBSI Sunwoda Battery Dataset
Shengyu, T. (2024). TBSI Sunwoda Battery Dataset [Data set]. Zenodo. https://doi.org/10.5281/zenodo.10715209 Electrical component
@@
Signals: current, voltage Prognosis Link

### v3

@@ 97 UCI - Condition monitoring of hydraulic systems Data Set
ZeMA - Center for Mechatronics and Automation Technology Saarbrücken, Germany
Mechatronic system
@@
Signals: efficiency, flow rate, power, pressure, temperature, vibration
Diagnosis Link
@@ 98 UCI - Mechanical Analysis 1990
F. Bergadano, A. Giordana, L. Saitta, (1990). University of Torino, Italy
Mechatronic system
@@
Signals: speed
Diagnosis Link
@@ 99 UCI - Robot Execution Failures Data Set
Universidade Nova de Lisboa, Monte da Caparica, Portugal
Robotic
@@
Signals: force, torque
Diagnosis Link
@@
Semeion Research Center of Sciences of Communication - Rome, Italy
Materials
@@
Signals: geometry, inspection data, operating condition
Diagnosis Link
@@
Lulea University of Technology: Martin del Campo Barraza, Sergio (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Sandin, Fredrik (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Strömbergsson, Daniel (Department of Engineering Sciences and Mathematics, Machine Elements)
Mechanical component
@@
Signals: vibration
Fault detection Link
@@
University Paderborn Kat
Mechanical component
@@
Signals: speed, temperature, torque
Diagnosis Link
@@
Virkler, D. A., Hillberry, B. M. and Goel, P. K. (1979). The Statistical Nature of Fatigue Crack Propagation. Journal of Engineering Materials and Technology, vol. 101, 148–153.
Materials
@@
Signals: crack length
Prognosis Link
@@
De Bruijn, B., Nguyen, T. A., Bucur, D., & Tei, K. (2016). Benchmark datasets for fault detection and classification in sensor data. In A. Ahrens, O. Postolache, & C. Benavente-Peces (Eds.), SENSORNETS 2016 - Proceedings of the 5th International Conference on Sensor Networks, pp. 185-195. SciTePress. https://doi.org/10.5220/0005637901850195
Electrical component
@@
Signals: light, temperature
Diagnosis Link
@@
Aimiyekagbon, O. K. (2024). Run-to-failure data set of ball bearings subjected to time-varying load and speed conditions. Zenodo, doi: 10.5281/zenodo.10805043, University Paderborn
Mechanical component
@@
Signals: temperature, vibration
Prognosis Link
@@
Jiangong Zhu, Yixiu Wang, Yuan Huang, R. Bhushan Gopaluni, Yankai Cao, Michael Heere, Martin J. Mühlbauer, Liuda Mereacre, Haifeng Dai, Xinhua Liu, Anatoliy Senyshyn, Xuezhe Wei, Michael Knapp, & Helmut Ehrenberg. (2022). Data-driven capacity estimation of commercial lithium-ion batteries from voltage relaxation [Data set]. Zenodo. https://doi.org/10.5281/zenodo.6405084
Electrical component
@@
Signals: current, temperature, voltage
Prognosis Link
@@
Zhang, Y., Tang, Q., Zhang, Y., Wang, J., Stimming, U., & Lee, A. A. (2020). Identifying degradation patterns of lithium ion batteries from impedance spectroscopy using machine learning [Data set]. Zenodo. https://doi.org/10.5281/zenodo.3633835
Electrical component
@@
Signals: current, temperature, voltage
Prognosis Link
108 Zenodo - Monitoring data railway bridge Leuven
Kristof Maes, & Geert Lombaert. (2020). Monitoring data for railway bridge KW51 in Leuven, Belgium, before, during, and after retrofitting [Data set]. Zenodo. https://doi.org/10.5281/zenodo.3745914
Building
Railway bridge
Signals: acceleration, displacement, strain
Diagnosis Link
@@
Huawei Munich Research Center
Mechatronic system
@@
Signals: humidity, vibration
Prognosis Link
@@
Shengyu, T. (2024). TBSI Sunwoda Battery Dataset [Data set]. Zenodo. https://doi.org/10.5281/zenodo.10715209
Electrical component
@@
Signals: current, voltage
Prognosis Link
Appendix B: Classification of Publicly Available Data Sets
[…]
Domain Application FD D HA P Sum across domain
Building Building 1 2
Bridge 1
Drive technology Aircraft engine* 5 12
Electric motor 1 1 1
Diesel engine 1
Helicopter engine 1
Propulsion system* 1 1
Electrical component Battery 1 15 25
Capacitor 2
Circuit breaker 1
Fuel cell 2
Sensor 1 1
Transistor 2
Manufacturing process Planarization system 1 8
Drilling 1
Electrophoresis painting 1
Milling 2 2
Turning 1
Material Aluminum plate 2 4
Polymer composite 1
Steel plate 1
Mechanical component Anemometer 1 26
Ball screw drive 1 1
[…]
Bogie 1
Gear 1 3
Shaft 1
Mechatronic system Air compressor 1 8
Air pressure system 1
Electromechanical ball screw* 1
Electromechanical device 1
Elevator 1
Hydraulic system 1
Rock drill 1
Validation door 1
Process technology Filtration 5 7
Pump 1 1
Production system Ion mill etching tool 1 8
Log data 1
Production line 2 3
Shrink-wrapper 1
Robotic Articulated robot 4 5
Linear robot 1
Unknown 2 2 1 5
all 18 35 3 54 110
Appendix C: Signals within Publicly Available Data Sets
Table 2: Signals used in at least two of the given data sets. Sorted by the sum of signal occurrences across the tasks of fault detection (FD), diagnosis (D), health assessment (HA), and prognosis (P).
Signal FD D HA P Sum across tasks
Vibration 4 16 0 14 34
Current 2 9 0 22 33
Temperature 4 8 0 21 33
Voltage 2 4 0 19 25
Speed 4 11 0 5 20
Pressure 0 5 1 12 18
Anonymized(1) 3 1 2 4 10
Flow Rate 0 1 0 9 10
Position 2 4 0 2 8
Torque 2 5 0 1 8
Operating Condition(2) 0 1 0 6 7
Acoustic Emission 0 2 0 3 5
Force 0 3 0 1 4
Acceleration(3) 1 2 0 0 3
Unknown(4) 2 0 0 1 3
Capacity 0 0 0 2 2
Humidity 0 1 0 1 2
Image 1 0 0 1 2
Impedance 0 0 0 2 2
Inspection Data 0 2 0 0 2
Power 1 1 0 0 2
Strain 0 1 0 1 2
(1): signal values changed/transformed,  (2): varying operating conditions during the life
(3): in terms of movement, not vibration, (4): signal (or sensor) type unknown